A +40.0 𝑛𝐢 charge is placed at the origin of a coordinate system, and a βˆ’30.0 𝑛𝐢 charge is placed in the π‘₯𝑦-plane at the point(π‘₯=0.200 π‘š,𝑦=βˆ’0.100 π‘š).

a) Determine the magnitude and the direction of the force that the positive charge exerts on the negative charge. Use π‘–π‘—π‘˜ notation for your result.

b) Determine the magnitude and direction of the force of the negative charge on the positive charge.


1
Expert's answer
2021-07-05T08:44:42-0400

(a)


"F=k\frac{q_1q_2}{r^2}=9\cdot10^9\cdot\frac{40\cdot10^{-9}\cdot30\cdot10^{-9}}{(\sqrt{0.2^2+0.1^2})^2}=0.000216\ (N)"


"\tan\alpha=-\frac{0.1}{0.2}=-0.5\to \alpha=153.435°" (relative to the positive direction of the x-axis)


"F_x=0.000216\cdot\cos153.435°=-1.93\cdot10^{-4}\ (N)"


"F_y=0.000216\cdot\sin153.435°=0.96\cdot10^{-4}\ (N)"


"\vec F_{p-n}=-1.93\cdot10^{-4}\ \vec i+0.96\cdot10^{-4}\ \vec j"


(b)


"\vec F_{n-p}=-\vec F_{p-n}"


"\tan\alpha=-\frac{0.1}{0.2}=-0.5\to \alpha=-26.56°" (relative to the positive direction of the x-axis)


"\vec F_{n-p}=1.93\cdot10^{-4}\ \vec i-0.96\cdot10^{-4}\ \vec j"